OpenStudy (anonymous):

Joe can mow the grass in 2 hr. Tom does the same job in 3 hr. How long would it take the two of them, working together, to mow the grass?

OpenStudy (anonymous):

3hrs+2hrs=5hrs 5hrs/2=2.5hrs the answer is 2.5 hrs

OpenStudy (anonymous):

^ Sorry, but that is very illogical thinking and is not the way to answer the question.

OpenStudy (anonymous):

sorry but the answer was right tho

OpenStudy (anonymous):

but please show how to do it correctly

OpenStudy (anonymous):

g = 1 (1 lawn of grass) g = 1/2t (Joe) + 1/3t(Tom) g = 3/6t + 2/6t g = 5/6t 1 = 5/6t t = 1.2 Correct answer is 1.2 hours. Let's call @Mertsj to check my work.

OpenStudy (anonymous):

ok dude or gal no problamo.

OpenStudy (anonymous):

also check my work please

OpenStudy (anonymous):

Let's call @Hero , since apparently @Mertsj isn't here.

OpenStudy (mertsj):

Joe's rate is 1/2 lawn per hour Tom's rate is 1/3 lawn per hour In x hours Joe mows x/2 lawns In x hours Tom mows x/3 lawns Together they mow 1 lawn So: \[\frac{x}{2}=\frac{x}{3}=1\]

OpenStudy (anonymous):

so we were both wrong

OpenStudy (mertsj):

Put a plus sign in there. Sorry

OpenStudy (anonymous):

Mertsj, the only possible value of x that would satisfy that equation is 0, which is obviously wrong.

hero (hero):

J = How many hours Joe can mow alone T = How many hours Tom can mow alone x = How many hours they can mow together \[\frac{J \times T}{J + T} = x\]

OpenStudy (anonymous):

Okay, never mind.

OpenStudy (mertsj):

\[\frac{x}{2}+ \frac{x}{3}=1\]

hero (hero):

\[\frac{2 \times 3}{2 + 3} = x\]

OpenStudy (anonymous):

So, x = 1.2, which is what I said before. Thanks Mertsj :).

OpenStudy (mertsj):

5x=6 x=6/5 hours to mow the lawn if both boys are working.

OpenStudy (anonymous):

losers

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

I think Mertsj and Hero deserve the best response because they explained it well, while I didn't really explain how I got 1.2 hours.